# Rotated Multi-Scale Interaction Network for Referring Remote Sensing Image Segmentation

### v2

@@ 5.4.1 Quantitative Results
To provide a comprehensive understanding of our designed model, we conduct a qualitative comparison of prediction results with the baseline. As depicted in Fig. 6, our model exhibits a remarkable capacity to precisely identify targets appearing at various scales based on expressions. Furthermore, it adeptly localizes tiny-scale objects within noisy backgrounds and robustly predicts objects appearing at different angles. In contrast, the predicted masks generated by the baseline model exhibit deficiencies, including missing portions and noticeable shifts.
@@ 6 Conclusion
In this paper, we introduce a Rotated Multi-Scale Interaction Network (RMSIN), a novel solution addressing the complex spatial scales and orientations found in RRSIS. The introduction of the Intra-scale Interaction Module and the Cross-scale Interaction Module in RMSIN specifically addresses the challenge of varied spatial scales found in aerial imagery. Furthermore, the integration of the Adaptive Rotated Convolution within RMSIN provides a robust solution for effectively handling the diverse orientations characteristic of such imagery. The extensive validation on our newly developed, comprehensive RRSIS-D dataset not only demonstrates RMSIN’s superior performance but also sets a new benchmark for future research.

### v3

@@ 5.4.1 Quantitative Results
We qualitatively compare our model with the baseline to provide a comprehensive understanding. As shown in Fig. 6, our model excels at identifying targets across various scales accurately, even within noisy backgrounds and at different angles. In contrast, the baseline model exhibits shortcomings such as missing parts and noticeable shifts in predicted masks.
@@ 6 Conclusion
In this paper, we propose RMSIN, a novel method adept at navigating the complex scales and orientations prevalent in RRSIS. By integrating the IIM and CIM, RMSIN capably addresses the wide range of spatial scales encountered in aerial imagery. Additionally, the implementation of the ARC offers a solid strategy for tackling various orientations. The construction of our expansive RRSIS-D dataset, featuring 17,402 image-caption-mask triplets provides an unparalleled resource in terms of scale and variety for rigorous evaluation. The comprehensive validation on the RRSIS-D dataset not only underscores RMSIN’s superior performance but also establishes a new benchmark for future research in this domain.
Acknowledgement
This work was supported by National Key R&D Program of China (No.2023YFB4502804), the National Science Fund for Distinguished Young Scholars (No.62025603), the National Natural Science Foundation of China (No. U21B2037, No. U22B2051, No. 62072389), the National Natural Science Fund for Young Scholars of China (No. 62302411), China Postdoctoral Science Foundation (No. 2023M732948), the Natural Science Foundation of Fujian Province of China (No.2021J01002, No.2022J06001), and partially sponsored by CCF-NetEase ThunderFire Innovation Research Funding (NO. CCF-Netease 202301).